# Simulation of Polynomial-Time Nondeterministic Turing Machines via Universal Nondeterministic Turing Machines

### v16

@@ 1.3. Our Approach
As a novel idea and attempt, we will apply simulation techniques. To the best of our knowledge, if we want to create a language not in a specific complexity class $\mathcal{C}$, the most obvious way is to apply the diagonalization technique or the lazy-diagonalization technique, because the simulation and diagonalization technique [Tur37, HS65] is a standard method (currently the only way) to prove lower bounds on uniform computing models (i.e., the Turing machine model); but, by the relation between the language $L(M)$ accepted by polynomial-time nondeterministic Turing machine $M$ and $\overline{L}(M)$ accepted by ${\rm co}\mathcal{NP}$ machine $M$, we will apply only the simulation technique. So, we first try to show that all polynomial-time nondeterministic Turing machines are enumerable. Then, we use a universal nondeterministic Turing machine to simulate all polynomial-time nondeterministic Turing machines, creating a language that is not in ${\rm co}\mathcal{NP}$. To stress it is “novel,” it is because this method is the first appearing in this paper, and we have never seen such an attempt appeared in previous works or such an idea in the relevant standard textbooks, such as [AHU74, AB09, Sip13, Pap94].
@@ 3. All Polynomial-time Nondeterministic Turing Machines Are Enumerable
In fact, by Definition 2.4, the above enumeration $e$ also gives an enumeration of all ${\rm co}\mathcal{NP}$ machines. To see this, we just force the non-trivial polynomial-time nondeterministic Turing machine in the enumeration $e$ to work under the condition of “there exists no accepting path.” Then, it is easy to see that each polynomial-time nondeterministic Turing machine corresponds to a ${\rm co}\mathcal{NP}$ machine, and vice versa.
@@ 5. Proving That $L_{s}\in\mathcal{NP}$
By Theorem 5.2, the nondeterministic Turing machine $U$ is also in the enumeration $e$. Or when working under the “there exists no accepting path” condition, the ${\rm co}\mathcal{NP}$ machine $U$ is also in the enumeration $e$, but $L(U)\neq\overline{L}(U)$. Indeed, since for any input $w\in\{0,1\}^{*}$, $w\in L(U)$ if and only if $w\not\in\overline{L}(U)$. Note that the universal nondeterministic Turing machine $U$ can simulate itself.
@@ 6. Breaking the Relativization Barrier
To explore what is behind this kind of mystery, let us dive into two cases: one is $\mathcal{P}^{O}=\mathcal{NP}^{O}={\rm co}\mathcal{NP}^{O}$, and another is $\mathcal{P}^{A}\neq\mathcal{NP}^{A}={\rm co}\mathcal{NP}^{A}$.
Note that Baker, Gill, and Solovay [BGS75]’s result also implies that for the same oracle $O$,
@@
First, note that $\mathcal{P}^{O}=\mathcal{NP}^{O}$, thus we have ${\rm co}\mathcal{P}^{O}={\rm co}\mathcal{NP}^{O}$. Clearly, in this scenario, to separate $\mathcal{NP}^{O}$ from ${\rm co}\mathcal{NP}^{O}$ is equivalent to separating $\mathcal{P}^{O}$ from ${\rm co}\mathcal{P}^{O}$. Moreover, since the set of all ${\rm co}\mathcal{P}^{O}$ machines is enumerable, the rest of the proof is similar to the Remark 4.2. Namely, the simulation technique is unable to create a language not in ${\rm co}\mathcal{P}^{O}$ via a universal deterministic oracle Turing machine $U^{O}$, because a polynomial-time deterministic oracle Turing machine under the “there exists no accepting path” condition is still a polynomial-time deterministic oracle Turing machine. In short, the “there exists at least one accepting path” condition and the “there exists no accepting path” condition for deterministic oracle Turing machines are symmetrical. Specifically, for an arbitrary deterministic oracle Turing machine $M^{O}$, if $M^{O}$ accepts the input $w$, then there is only one accepting path, and if $M^{O}$ rejects the input $w$, then there is also only one rejecting path. This is different from nondeterministic oracle Turing machines. If we apply the same techniques, i.e., the simulation techniques, to create a language not in ${\rm co}\mathcal{P}^{O}$, then the universal deterministic oracle Turing machine $U^{O}$ (it runs within time $O(n^{k})$ for any $k\in\mathbb{N}_{1}$) needed to be constructed accepts the language $L(U^{O})\in\mathcal{P}^{O}$ (the proof is similar to that of Theorem 6.4 below). But, by the symmetrical properties between the “there exists at least one accepting path” condition and the “there exists no accepting path” condition with respect to deterministic oracle Turing machines, we clearly have that $L(U^{O})\in{\rm co}\mathcal{P}^{O}$. Thus, we are unable to claim that $L(U^{O})\notin{\rm co}\mathcal{P}^{O}$.

### v17

@@ 1.3. Our Approach
As a novel idea and attempt, we will apply simulation techniques. To the best of our knowledge, if we want to create a language not in a specific complexity class $\mathcal{C}$, the most obvious way is to apply the diagonalization technique or the lazy-diagonalization technique, because the simulation and diagonalization technique [Tur37, HS65] is a standard method (currently the only way) to prove lower bounds on uniform computing models (i.e., the Turing machine model). But, we find that it is hard to apply standard diagonalization or lazy-diagonalization techniques to produce a language that not in ${\rm co}\mathcal{NP}$. Instead, by the relation between the language $L(M)$ accepted by polynomial-time nondeterministic Turing machine $M$ and $\overline{L}(M)$ accepted by ${\rm co}\mathcal{NP}$ machine $M$, we will apply only simulation techniques. So, we first try to show that all polynomial-time nondeterministic Turing machines are enumerable. Then, we use a universal nondeterministic Turing machine to simulate all polynomial-time nondeterministic Turing machines, creating a language that is not in ${\rm co}\mathcal{NP}$. To stress it is “novel,” it is because this method is the first appearing in this paper, and we have never seen such an attempt appeared in previous works or such an idea in the relevant standard textbooks, such as [AHU74, AB09, Sip13, Pap94].
@@ 3. All Polynomial-time Nondeterministic Turing Machines Are Enumerable
In fact, by Definition 2.4, the above enumeration $e$ also gives an enumeration of all ${\rm co}\mathcal{NP}$ machines. To see this, we just force the non-trivial polynomial-time nondeterministic Turing machine in the enumeration $e$ to work in the condition of “there exists no accepting path.” Then, it is easy to see that each polynomial-time nondeterministic Turing machine corresponds to a ${\rm co}\mathcal{NP}$ machine, and vice versa.
@@ 5. Proving That $L_{s}\in\mathcal{NP}$
By Lemma 2.7 and Theorem 5.2, the single-tape nondeterministic Turing machine $U_{1}$ that is equivalent to the nondeterministic Turing machine $U$ (i.e., $L(U_{1})=L(U)$) is also in the enumeration $e$. Or when working in the “there exists no accepting path” condition, the single-tape ${\rm co}\mathcal{NP}$ machine $U_{1}$ is also in the enumeration $e$. Specifically, we know from the proof of Theorem 5.2 that, for any $i\in\mathbb{N}_{1}$, we can make the multi-tape universal nondeterministic Turing machine $U$ run at most $|w|^{i+1}$ steps for all input $w\in\{0,1\}^{*}$ (i.e., $U$ runs within time $O(n^{i+1})$); by Lemma 2.7, such a machine is equivalent to a single-tape nondeterministic $O(n^{2(i+1)})$ time-bounded Turing machine $U_{i}$, i.e., $L_{s}^{i}=L(U_{i})$, and the single-tape nondeterministic $n^{2(i+1)}+2(i+1)$ time-bounded Turing machine $(U_{i},2(i+1))$ is in the enumeration $e$. But we have that $L_{s}\neq\overline{L}(U)$ (note that $L_{s}=L(U)$). Indeed, since for any input $w\in\{0,1\}^{*}$, $w\in L_{s}$ if and only if $w\not\in\overline{L}(U)$.
@@ 6. Breaking the Relativization Barrier
To explore what is behind this kind of mystery, let us dive into two cases: one is $\mathcal{P}^{O}=\mathcal{NP}^{O}={\rm co}\mathcal{NP}^{O}$ for some oracle $O$, and another is $\mathcal{P}^{A}\neq\mathcal{NP}^{A}={\rm co}\mathcal{NP}^{A}$ for some oracle $A$.
Note that, the result of Baker, Gill, and Solovay [BGS75] also implies that for the same oracle $O$,
@@
First, note that $\mathcal{P}^{O}=\mathcal{NP}^{O}$, thus we have ${\rm co}\mathcal{P}^{O}={\rm co}\mathcal{NP}^{O}$. Clearly, in this scenario, to separate $\mathcal{NP}^{O}$ from ${\rm co}\mathcal{NP}^{O}$ is equivalent to separating $\mathcal{P}^{O}$ from ${\rm co}\mathcal{P}^{O}$. Moreover, if the set of all ${\rm co}\mathcal{P}^{O}$ machines is not enumerable, then of course we can not apply simulation techniques. Thus, we assume that the set of ${\rm co}\mathcal{P}^{O}$ machines is enumerable, as the premise of the theorem states. Since we have assumed that the set of all ${\rm co}\mathcal{P}^{O}$ machines is enumerable, the rest of the proof is similar to the Remark 4.2. Namely, the simulation technique is unable to create a language not in ${\rm co}\mathcal{P}^{O}$ via a universal deterministic oracle Turing machine $U^{O}$, because a polynomial-time deterministic oracle Turing machine under the “there exists no accepting path” condition is still a polynomial-time deterministic oracle Turing machine. In short, the “there exists at least one accepting path” condition and the “there exists no accepting path” condition for deterministic oracle Turing machines are symmetrical. Specifically, for an arbitrary deterministic oracle Turing machine $M^{O}$, if $M^{O}$ accepts the input $w$, then there is only one accepting path, and if $M^{O}$ rejects the input $w$, then there is also only one rejecting path. This is different from nondeterministic oracle Turing machines. If we apply the same techniques, i.e., the simulation techniques, to create a language not in ${\rm co}\mathcal{P}^{O}$, then the universal deterministic oracle Turing machine $U^{O}$ (it runs within time $O(n^{k})$ for any $k\in\mathbb{N}_{1}$) needed to be constructed accepts the language $L(U^{O})\in\mathcal{P}^{O}$ (the proof is similar to that of Theorem 6.4 below). But, by the symmetrical properties between the “there exists at least one accepting path” condition and the “there exists no accepting path” condition with respect to deterministic oracle Turing machines, we clearly have that $L(U^{O})\in{\rm co}\mathcal{P}^{O}$. Thus, we are unable to claim that $L(U^{O})\notin{\rm co}\mathcal{P}^{O}$.